# RILe: Reinforced Imitation Learning

### v2

@@ 3.4 Adversarial Imitation Learning (AIL) and Adversarial Inverse Reinforcement Learning (AIRL)
Imitation Learning (IL) aims to directly approximate the expert policy from given expert trajectory samples $\tau_{E}$. It can be formulated as
$\text{IL}(\tau_{E})=\operatorname*{arg\,min}_{\pi}E_{(s,a)\sim\tau_{E}}[L(\pi(\cdot|s),a)],$ (3)
where $L$ is a loss function, that captures the difference between policy and expert data.
GAIL [3] introduces an adversarial imitation learning setting by quantifying the difference between the agent and the expert with a discriminator $D_{\phi}(s,a)$, parameterized by $\phi$ (Fig. 1-(c)).
The discriminator distinguishes between between expert-generated state-action pairs $(s,a)\sim\tau_{E}$ and non-expert state-action pairs $(s,a)\notin\tau_{E}$. The goal of GAIL is to find the optimal policy that fools the discriminator while maximizing an entropy constraint. The optimization is formulated as a zero-sum game between the discriminator $D_{\phi}(s,a)$ and the policy $\pi$:
$\min_{\pi}\max_{\phi}E_{\pi}[\text{log}\ D_{\phi}(s,a)]+E_{\tau_{E}}[\text{log}\ (1-D_{\phi}(s,a))]-\lambda H(\pi).$ (4)
@@
AIRL [6] extends this adversarial approach to inverse reinforcement learning, aiming to recover a reward function decoupled from environment dynamics (Fig. 1-(c)). The key insight of AIRL is to structure the discriminator as:
@@
where $f_{\phi}(s,a,s^{\prime})=r_{\psi}(s,a)+\gamma V_{\phi}(s^{\prime})-V_{\phi}(s)$. Here, $r_{\psi}(s,a)$ represents the learned reward function that is decoupled from the environment dynamics, $\gamma V_{\phi}(s^{\prime})-V_{\phi}(s)$. The AIRL optimization problem is formulated equivalently to GAIL (see Eqn. 4). The reward function $r_{\psi}(s,a)$ is learned through minimizing the cross-entropy loss inherent in this adversarial setup. Therefore, the reward function remains tightly coupled with the discriminator’s learning process.
@@ 4 RILe: Reinforced Imitation Learning
We propose Reinforced Imitation Learning (RILe) to learn the reward function and acquire a policy that emulates expert-like behavior simultaneously in one learning process. Our framework consists of three key components: a student agent, a trainer agent, and a discriminator (Fig. 2).
Intuitively, the student agent learns the action policy and the trainer agent learns the reward function. Both agents are trained simultaneously via RL with the help of an adversarial discriminator. Building on AIL, we utilize the discriminator to enable joint training of policy and reward function. Unlike traditional IRL, the discriminator provides immediate feedback to the trainer, eliminating the need to wait for the student agent’s convergence.
The trainer agent, trained via RL, is rewarded by the discriminator, and thus motivated to maximize the cumulative reward from the discriminator. The trainer learns to craft a reward function that promotes expert-like behavior in the student, even in complex scenarios. Crucially, the trainer’s RL-based approach optimizes for long-term performance rather than immediate feedback. This allows the trainer to encourage exploration and gradual improvement in three main ways. First, the trainer can associate its reward-giving actions with future improvements in the student’s behavior, even if those improvements take many steps to manifest. Second, the trainer can provide small positive rewards for actions that move the student in the right direction, even if they don’t immediately result in expert-like behavior. Finally, the trainer can adapt its reward function based on the student’s current capabilities, providing a learning path that gradually guides the student towards expert behavior. This approach enables RILe to overcome limitations of previous methods, particularly in complex tasks requiring extensive exploration, by promoting the discovery of expert-like strategies even when the student’s initial policy significantly diverges from expert behavior.
In the following, we define the components of RILe and explain how they can efficiently learn behavior from imperfect data.
@@
Figure 2: Framework overview. The framework consists of three key components: a student agent, a trainer agent, and a discriminator. The student agent learns a policy $\pi_{S}$ by interacting with an environment, and the trainer agent learns a reward function as a policy $\pi_{T}$. (1) The student receives the environment state $s^{S}$. (2) The student takes an action $a^{S}$, forwards it to the environment which is updated based on $a^{S}$. (3) The student forwards its state and action to the trainer, whose state is $s^{T}=(s^{S},a^{S})$. (4) Trainer, $\pi_{T}$, evaluates the state action pair of the student agent $s^{T}=(s^{S},a^{S})$ and chooses an action $a^{T}$ that then becomes the reward of the student agent $a^{T}=r^{S}$. (5) The trainer agent forwards the $s^{T}=(s^{S},a^{S})$ to the discriminator. (6) Discriminator compares student state-action pair with expert demonstrations ($s^{D}$). (7) Discriminator gives reward to the trainer, based on the similarity between student- and expert-behavior.
@@
The student agent aims to learn a policy $\pi_{S}$ by interacting with an environment in a standard RL setting within an MDP. For each of its actions $a^{S}\in A$, the environment returns a new state $s^{S}\in S$. However, rather than from a hand-crafted reward function, the student agent receives its reward from the policy of the trainer agent $\pi_{T}$. Therefore, the reward function is represented by the trainer policy. Thus, the student agent is guided by the actions of the trainer agent, i.e., the action of the trainer is the reward of the student: $r^{S}=\pi_{T}((s^{S},a^{S}))$. The optimization problem of the student agent is then defined as
$\min_{\pi_{S}}-E_{(s^{S},a^{S})\sim\pi_{S}}[\pi_{T}\left((s^{S},a^{S})\right)].$ (6)
@@
The discriminator differentiates between expert-generated state-action pairs $(s,a)\sim\tau_{E}$ and state-action pairs stemming from the policy of the student $(s,a)\sim\pi_{S}$. In RILe, the discriminator is defined as a feed-forward deep neural network, parameterized by $\phi$. Hence, the optimization problem is
$\max_{\phi}E_{(s,a)\sim\tau_{E}}[\text{log}(D_{\phi}(s,a))]+E_{(s,a)\sim\pi_{S}}[\text{log}(1-D_{\phi}(s,a))].$ (7)
To provide effective guidance, the discriminator needs to accurately distinguish whether a given state-action pair originates from the expert distribution $(s,a)\sim\tau_{E}$ or not $(s,a)\notin\tau_{E}$. The feasibility of this discrimination has been demonstrated by GAIL [3]. The according lemma and proof are presented in the Appendix A.
@@
The trainer agent aims to guide the student to approximate expert behavior by operating as its reward mechanism. Since the trainer agent takes the role of a reward function for the student, a new MDP is defined for the trainer agent: $\text{MDP}_{\text{T}}:(S_{T},A_{T},R_{T},T_{T},K,\gamma)$, where $S_{T}:S\ \text{x}\ A$ is the state space that consists all possible state action pairs from the standard MDP, defined in 3.1. $A_{T}$ is the action space, a mapping from $S_{T}\ \rightarrow\ \mathbb{R}$, so the action is a scalar value. $R_{T}(s,a)$ is the reward function where $s\in S_{T}$ and $a\in A_{T}$. $T_{T}=\{P(\cdot|s,a)\}$ is the transition dynamics where $P(\cdot|s,a)$ is defined as the state distribution upon taking action $a\in A_{T}$ in state $s\in S_{T}$. $K_{T}$ is the initial state distribution.
The trainer agent learns a policy $\pi_{T}$ that produces adequate reward signals to guide the student agent, by learning in a standard RL setting, within $MDP_{T}$. Since the state space of $MDP_{T}$ is defined over state-action pairs of $MDP_{S}$, the state of the trainer comprises the state-action pair of the student $s^{T}=(s^{S},a^{S})\in S_{T}$. It generates a scalar action $a^{T}$ which is given to the student agent as reward $r^{S}$, and bounded between -1 to 1. To help the trainer understand how its actions impact the reward it receives, we define the reward function such that it multiplies the scaled discriminator’s output by trainer’s actions. Therefore, the trainer agent’s reward function is defined as $R^{T}=\upsilon(D_{\phi}(s^{T}))(a^{T})$, where $D_{\phi}(s^{T})$ is the output of the discriminator and $\upsilon(x)=2x-1$ is the scaling function. By incorporating $a^{T}$ into the reward function, the trainer learns to adjust its policy based on the effectiveness of its previous actions.
The optimization problem of the trainer can be defined as
$\max_{\pi_{T}}E_{\begin{subarray}{c}(s,a)\sim\pi_{S}\\ a^{T}\sim\pi_{T}\end{subarray}}[\upsilon(D_{\phi}(s,a))a^{T}].$ (8)
@@
RILe combines the three components defined previously in order to find a student policy that mimics expert behaviors presented in $\tau_{E}$. In RILe, the student policy $\pi_{S}$ and the trainer policy $\pi_{T}$ can be trained via any single-agent online reinforcement learning method. The training algorithm is given in Appendix D. Overall, the student agent aims to recover the optimal policy $\pi_{S}^{*}$ defined as
$\pi_{S}^{*}=\operatorname*{arg\,max}_{\pi_{S}}E_{(s^{S},a^{S})\sim\pi_{S}}\left[\sum_{t=0}^{\infty}\gamma^{t}[\pi_{T}\left((s^{S}_{t},a^{S}_{t})\right)\right].$ (9)
At the same time, the trainer agent aims to recover the optimal policy $\pi_{T}^{*}$ as
$\pi_{T}^{*}=\operatorname*{arg\,max}_{\pi_{T}}E_{\begin{subarray}{c}s^{T}\sim\pi_{S}\\ a^{T}\sim\pi_{T}\end{subarray}}\left[\sum_{t=0}^{\infty}\gamma^{t}[\upsilon(D_{\phi}(s^{T}_{t}))a^{T}\right].$ (10)
Lemma 1: Given the discriminator $D_{\phi}$, the trainer agent optimizes its policy $\pi^{\theta_{T}}$ via policy gradients to provide rewards that guide the student agent to match expert’s state-action distributions.
Proof for Lemma 1 The assumptions are presented in Appendix A. The expert’s state-action distribution is denoted by $p_{\text{expert}}(s,a)$. The role of the trainer is to provide a reward signal to the student that encourages the approximation of $p_{\text{expert}}(s,a)$ as closely as possible.
We have $D_{\phi}:\mathcal{S}\times\mathcal{A}\rightarrow[0,1]$ as the discriminator, parameterized by $\phi$, which outputs the likelihood that a given state-action pair $(s,a)$ originates from the expert, as opposed to the student. The trainer’s policy $\pi^{\theta_{T}}$, parameterized by $\theta_{T}$, aims to maximize the likelihood under the discriminator’s assessment, thus encouraging the student agent to generate state-action pairs drawn from a distribution resembling $p_{\text{expert}}(s,a)$.
The Value and Q functions of the trainer, conditioned on the rewards provided by the discriminator, are defined in terms of expected cumulative discriminator rewards. The value function for the trainer’s policy parameters $\theta_{T}$ at state $s_{t}$, simplified from $s^{T}_{t}$ for better readability, is given by:
$V^{\theta_{T}}(s_{t})=\mathbb{E}_{\pi^{\theta_{T}}}\left[\sum_{k=0}^{\infty}\gamma^{k}r_{t+k}\mid s_{t}\right],$ (11)
where $r_{t+k}=\upsilon(D_{\phi}(s_{t+k}))a_{t+k}$. Similarly, the Q-function for taking action $a_{t}$ in state $s_{t}$ and then following policy $\pi^{\theta_{T}}$ can be written as:
$Q^{\theta_{T}}(s_{t},a_{t})=D_{\phi}(s_{t})+\gamma\mathbb{E}_{\pi^{\theta_{T}}}\left[V^{\theta_{T}}(s_{t+1})\mid s_{t},a_{t}\right].$ (12)
The trainer’s policy optimization is done by maximizing the following clipped surrogate objective function:
$\begin{split}L^{CLIP}(\theta_{T})=&\mathbb{E}_{(s_{t},a_{t})\sim\pi^{\theta_{T}}}\Big{[}\min\Big{(}\frac{\pi^{\theta_{T}}(a_{t}|s_{t})}{\pi^{\theta_{T_{old}}}(a_{t}|s_{t})}A^{\theta_{T_{old}}}(s_{t},a_{t}),\\ &\qquad\qquad\qquad\text{clip}(\frac{\pi^{\theta_{T}}(a_{t}|s_{t})}{\pi^{\theta_{T_{old}}}(a_{t}|s_{t})},1-\epsilon,1+\epsilon)A^{\theta_{T_{old}}}(s_{t},a_{t})\Big{)}\Big{]},\end{split}$ (13)
where $A^{\theta_{T_{old}}}(s_{t},a_{t})$ is the advantage function computed as $Q^{\theta_{T_{old}}}(s_{t},a_{t})-V^{\theta_{T_{old}}}(s_{t})$, defined with respect to the trainer’s old policy parameters $\theta_{T_{old}}$.
By expressing the advantage using the reward from the discriminator, we explicitly tie the policy gradient updates to the discriminator’s output, emphasizing the shaping of $\pi^{\theta_{S}}$ to match $p_{\text{expert}}(s,a)$:
$A^{\theta_{T_{old}}}(s_{t},a_{t})=D_{\phi}(s_{t})+\gamma\mathbb{E}_{\pi^{\theta_{S}}}\left[V^{\theta_{T_{old}}}(s_{t+1})\right]-V^{\theta_{T_{old}}}(s_{t}).$ (14)
During each policy update, the objective in Equation 13 is maximized, driving parameter updates to favor actions that elicit higher rewards from the discriminator – effectively the actions that better align with expert behavior:
$\theta_{T}\leftarrow\theta_{T}+\alpha_{\pi}\nabla_{\theta_{T}}L^{CLIP}(\theta_{T}).$ (15)
The update rule in Equation 15, driven by cumulative rewards from the discriminator, incrementally adapts the trainer’s policy to reinforce student behaviors that are indistinguishable from those of the expert according to $D_{\phi}$. The trainer’s policy is guided to facilitate the student’s approximation of $p_{\text{expert}}(s,a)$, thereby fulfilling its role in the imitation learning process.
@@ 5 Experiments
We evaluate the performance of RILe by addressing three key questions:
@@
    How does RILe’s adaptive reward function evolve compared to baseline methods and how does this evolution enhance the learning process?
@@
    Is RILe efficient and scalable to high-dimensional continuous control tasks?
@@
    Can RILe use expert-data explicitly to imitate expert behavior?
To answer the first question, we compare RILe’s performance with AI(R)L baselines in a maze setting, where we demonstrate how the trainer agent modifies the reward function to guide the student during training. For the second question, we evaluate RILe’s effectiveness in imitating motion-capture data within continuous robotic control tasks, using LocoMujoco [20] as the environment. To answer the last question, we use a simulated humanoid character from MuJoCo [21, 22] to evaluate RILe’s performance when expert data is explicitly used by the agents. Additional experimental results are provided in the Appendix, where we evaluate the robustness of the learned reward function in changing environments and analyze the noise resilience of our method.
Baselines We compare RILe with four baseline methods: Behavioral cloning (BC [7, 23]), adversarial imitation learning (GAIL [3]), adversarial inverse reinforcement learning (AIRL [6]) and inverse reinforcement learning (IQ-Learn [24]).

### v3

@@ 3.4 Adversarial Imitation Learning (AIL) and Adversarial Inverse Reinforcement Learning (AIRL)
[…]
GAIL [5] introduces an adversarial imitation learning setting by quantifying the difference between the agent and the expert with a discriminator $D_{\phi}(s,a)$, parameterized by $\phi$ (Fig. 1-(c)). The discriminator distinguishes between between expert-generated state-action pairs $(s,a)\sim\tau_{E}$ and non-expert ones $(s,a)\notin\tau_{E}$. The goal of GAIL is to find the optimal policy that fools the discriminator while maximizing an entropy constraint. The optimization is formulated as a zero-sum game between the discriminator $D_{\phi}(s,a)$ and the policy $\pi$:
$\min_{\pi}\max_{\phi}\mathbb{E}_{\pi}[\text{log}\ D_{\phi}(s,a)]+\mathbb{E}_{\tau_{E}}[\text{log}\ (1-D_{\phi}(s,a))]-\lambda H(\pi).$ (3)
@@
AIRL [8] extends AIL to inverse reinforcement learning, aiming to recover a reward function decoupled from environment dynamics (Fig. 1-(c)). AIRL structures the discriminator as:
@@
where $f_{\phi}(s,a,s^{\prime})=r_{\psi}(s,a)+\gamma V_{\phi}(s^{\prime})-V_{\phi}(s)$. Here, $r_{\psi}(s,a)$ represents the learned reward function that is decoupled from the environment dynamics, $\gamma V_{\phi}(s^{\prime})-V_{\phi}(s)$. The AIRL optimization problem is formulated equivalently to GAIL (see Eqn. 3). The reward function $r_{\psi}(s,a)$ is learned through minimizing the cross-entropy loss inherent in this adversarial setup. Therefore, the reward function remains tightly coupled with the discriminator’s learning process.
@@ 4 RILe: Reinforced Imitation Learning
We propose Reinforced Imitation Learning (RILe) to jointly learn a reward function and a policy that emulates expert-like behavior within a single learning process. RILe introduces a novel trainer–student dynamic, as illustrated in Figure 2.
In RILe, the student agent learns an action policy by interacting with the environment, while the trainer agent learns a reward function that effectively guides the student toward expert-like behavior. Both agents are trained simultaneously via reinforcement learning, with an assistance from an adversarial discriminator. Specifically, the trainer queries the discriminator, which judges how expert-like the student’s behavior is, and then optimizes the reward function based on that feedback on-the-fly. Unlike traditional AIL, where the discriminator effectively is employed as the reward function for the student, RILe introduces a trainer agent to provide fine-grained feedback to the student, while avoiding IRL’s iterative computational expense.
The trainer agent plays the key role in RILe. Trained via RL, the trainer explores different reward designs and learns to provide gradually tailored feedback to the student by maximizing the cumulative rewards it receives from the discriminator. This approach equips RILe with two key advantages over existing IRL/AIRL/AIL frameworks: (1) On-the-fly reward function learning via RL: The reward function is learned continuously with RL, enabling the trainer to explore different reward options and account for long-horizon effects of its signals, (2) Context-sensitive guidance: The trainer adjusts its reward outputs in response to the student’s current policy, thereby encouraging the student to explore suboptimal actions that ultimately guide it closer to expert behavior. By providing tailored feedback at different stages of training, RILe addresses the limitations of prior methods, particularly in high-dimensional tasks.
In the remainder of this section, we define the components of RILe and explain how they jointly learn from expert demonstrations.
@@
Figure 2: Reinforced Imitation Learning (RILe). The framework consists of three key components: a student agent, a trainer agent, and a discriminator. The student agent learns a policy $\pi_{S}$ by interacting with an environment, and the trainer agent learns a reward function as a policy $\pi_{T}$. (1) The student receives the environment state $s^{S}$. (2) The student takes an action $a^{S}$, forwards it to the environment which is updated based on $a^{S}$. (3) The student forwards its state and action to the trainer, whose state is $s^{T}=(s^{S},a^{S})$. (4) Trainer, $\pi_{T}$, evaluates the state action pair of the student agent $s^{T}=(s^{S},a^{S})$ and chooses an action $a^{T}$ that then becomes the reward of the student agent $a^{T}=r^{S}$. (5) The trainer agent forwards the $s^{T}=(s^{S},a^{S})$ to the discriminator. (6) Discriminator compares student state-action pair with expert demonstrations ($s^{D}$). (7) Discriminator gives reward to the trainer, based on the similarity between student- and expert-behavior.
@@
The student agent learns a policy $\pi_{S}$ by interacting with an environment in a standard RL setting within an MDP. For each of its actions $a^{S}\in A$, the environment returns a new state $s^{S}\in S$. However, instead of using a handcrafted reward function, the student’s reward comes from the trainer agent’s policy, $\pi_{T}$. Therefore, the reward function is represented by the trainer policy. Thus, the student agent is guided by the actions of the trainer agent, i.e., the action of the trainer is the reward of the student: $r^{S}=\pi_{T}((s^{S},a^{S}))$. The optimization problem of the student agent is then defined as
$\min_{\pi_{S}}-\mathbb{E}_{(s^{S},a^{S})\sim\pi_{S}}[\pi_{T}\left((s^{S},a^{S})\right)].$ (5)
@@
The discriminator differentiates between expert-generated state-action pairs, $(s,a)\sim\tau_{E}$, and pairs from the student, $(s,a)\sim\pi_{S}$. In RILe, the discriminator is defined as a feed-forward deep neural network, parameterized by $\phi$. Its objective is:
$\max_{\phi}\mathbb{E}_{(s,a)\sim\tau_{E}}[\text{log}(D_{\phi}(s,a))]+\mathbb{E}_{(s,a)\sim\pi_{S}}[\text{log}(1-D_{\phi}(s,a))].$ (6)
To provide effective guidance, the discriminator must accurately identify whether a given state–action pair originates from the expert distribution $(s,a)\sim\tau_{E}$ or not $(s,a)\notin\tau_{E}$. GAIL [5] established the feasibility of such a discriminator (see Appendix B for details).
@@
The trainer agent guides the student toward expert behavior by serving as its reward mechanism. Since the trainer does not directly observe the student’s policy $\pi_{S}$, we model the trainer’s environment as a Partially Observable MDP (POMDP): $\text{POMDP}_{T}=(S_{T},A_{T},\Omega_{T},T_{T},O_{T},R_{T},\gamma)$. The state space $S_{T}=S\times A\times\pi_{S}$ includes all possible state-action pairs from the standard MDP and the student’s policy $\pi_{S}$, which is hidden from the trainer, introducing partial observability. The trainer’s action space, $A_{T}$, consists of scalar values. Formally, $A_{T}$ is defined as a mapping from $S_{T}\ \rightarrow\ \mathbb{R}$. The observation space $\Omega_{T}=S\times A$ consists of the observable state-action pairs of the student. The transition dynamics $T_{T}$ and the observation function $O_{T}$ are defined formally in Appendix A. The reward function $R_{T}(s^{T},a^{T})$ evaluates the effectiveness of the trainer’s action in guiding the student, where $s^{T}=(s^{S},a^{S})$ is the observation of the trainer. $\gamma$ is the discount factor.
Within this POMDP, the trainer learns a policy $\pi_{T}$ that produces helpful reward signals for $\pi_{S}$. The trainer observes only the student’s state–action pair, $s^{T}=(s^{S},a^{S})\in S\times A$, not $\pi_{S}$ itself. It then outputs a scalar action $a^{T}\in[-1,1]$, which is provided to the student as the reward, $r^{S}$.
If the trainer’s reward depends only on the discriminator’s output, the trainer receives the same reward regardless of whether it rewards or penalizes the student, yielding no immediate feedback on its choices. For instance, if the student behaves like the expert and the discriminator outputs $\approx$ 1, the trainer should ideally reward the student (action, $a^{T}$, $\approx$ 1). But if the trainer’s action is not factored into its own reward, it gains no immediate signal whether rewarding or punishing the student was effective, since it receives the same reward in either case. This ambiguity forces extensive trial and error. To address this, we define the trainer reward as:
$R^{T}=\upsilon(D_{\phi}(s^{T}))(a^{T})$ (7)
where $\upsilon(x)=2x-1$ scales the discriminator’s output, making it symmetric around zero. Including $a^{T}$ in the trainer’s reward ensures the trainer effectively learns from its own actions. Formally, the trainer’s objective is:
$\max_{\pi_{T}}\mathbb{E}_{\begin{subarray}{c}(s,a)\sim\pi_{S}\\ a^{T}\sim\pi_{T}\end{subarray}}[\upsilon(D_{\phi}(s,a))a^{T}].$ (8)
@@
RILe brings together these three components, student, trainer, and discriminator, to discover a student policy that imitates expert behaviors in $\tau_{E}$. Both $\pi_{S}$ and $\pi_{T}$ can be trained via any single-agent RL method. The overall training algorithm is detailed in Appendix J.
The student agent aims to recover the optimal policy $\pi_{S}^{*}$:
$\pi_{S}^{*}=\operatorname*{arg\,max}_{\pi_{S}}\mathbb{E}_{(s^{S},a^{S})\sim\pi_{S}}\left[\sum_{t=0}^{\infty}\gamma^{t}[\pi_{T}\left((s^{S}_{t},a^{S}_{t})\right)]\right].$ (9)
Simultaneously, the trainer aims to recover $\pi_{T}^{*}$:
$\pi_{T}^{*}=\operatorname*{arg\,max}_{\pi_{T}}\mathbb{E}_{\begin{subarray}{c}s^{T}\sim\pi_{S}\\ a^{T}\sim\pi_{T}\end{subarray}}\left[\sum_{t=0}^{\infty}\gamma^{t}[\upsilon(D_{\phi}(s^{T}_{t}))a^{T}_{t}]\right].$ (10)
By optimizing these objectives together, RILe efficiently learns both a reward function and a policy in high-dimensional settings where traditional AIL or IRL methods often struggle. Details on specific training strategies are provided in Appendix C.
@@ 5 Experiments
We evaluate the performance of RILe by addressing four key questions:
@@
    How does RILe’s reward-learning strategy differs from AIL and AIRL?
@@
    How adaptive is RILe’s learned reward function?
@@
    How does RILe perform in both low- and high-dimensional continuous-control problem compared to AIL/AIRL/IRL?
  4. 4.
    How explicitly using expert-data within RILe’s training process affects the context-sensitive reward learning?
Baselines We compare RILe with seven baseline methods: Behavioral cloning (BC [4, 21], BCO [10]), adversarial imitation learning (GAIL [5], GAIfO [9] and DRAIL [11]), adversarial inverse reinforcement learning (AIRL [8]), and inverse reinforcement learning (IQ-Learn [7]). DRAIL [11] introduces a diffusion-based discriminator implementation, which is applied to both GAIL and RILe, and referred as DRAIL-GAIL and DRAIL-RILe.
Additional experimental details are provided in the Appendix D, and hyperparameter selections are discussed in the Appendix H.